how many pounds are in 1 1/4 tons
Found 2 solutions by jim_thompson5910, stanbon:
Answer by jim_thompson5910(35256)

1 ton = 2000 pounds

(5/4)*(1 ton) = (5/4)*(2000 pounds)

5/4 tons = 2500 pounds

1 1/4 tons = 2,500 pounds

Answer by stanbon(75887)

how many pounds are in 1 1/4 tons
----
Assuming 1 ton = 2000 lbs (american wt.),
1 1/4 tone = (5/4)2000 = 5*500 = 2500 lbs
=================
Cheers,
Stan H.